The sum of two numbers is 20 the sum of the smaller and 4 times the larger is 56 . Find the numbers
Answer by algebrahouse.com(1659)

x = smaller number
y = larger number

x + y = 20 {sum of two numbers is 20}
x + 4y = 56 {sum of smaller and 4 times larger is 56}
-----------
-3y = -36 {subtracted the two equations, eliminating x}
y = 12 {divided each side by -3}

x + y = 20 {first original equation}
x + 12 = 20 {substituted 12, in for x, into first original equation}
x = 8 {subtracted 12 from each side}

smaller number is 8
larger number is 12

For more help from me, visit: www.algebrahouse.com